OpenStudy (nina21):

What is another way to write -3 • (4 + 7) A -3 • 4 + 7 B -3 • 4 • 7 *** C -3 • 4 + 4 • 7 D -3 • 4 + (-3) • 7

OpenStudy (mathstudent55):

Here is the distributive property of multiplication over addition a(b + c) = ab + ac Examples: 2(5 + 7) = 2 * 5 + 2 * 7 -4(8 + 5) = -4 * 8 + (-4) * 5 Follow the rule and the examples and apply the property to your problem. What do you get?

OpenStudy (nina21):

D

OpenStudy (mathstudent55):

Correct.

OpenStudy (nina21):

thanks

OpenStudy (mathstudent55):

Here is a geometric explanation of the distributive law. 

OpenStudy (mathstudent55):

The large rectangle has length b + c. Its width is a. The area of the large rectangle is length times width, or a * (b + c), or simply a(b + c). The small rectangle to the left has length and width a and b and area ab. The small rectangle to the right has length and width a and c and area ac. The area of the large rectangle is equal to the sum of the areas of the two small rectangles, so a(b + c) = ab + ac

OpenStudy (mathstudent55):

You're welcome.